# On the use of the Gram matrix for multivariate functional principal components analysis

### v1

@@ Appendix B Derivation of the eigencomponents
Therefore, in order to have an orthonormal basis of eigenfunctions, we normalise the eigenfunctions $\phi_{k}$ from Equation (73) by $1/\sqrt{l_{k}}$. Concerning the estimation of the scores, for $n=1,\dots,N$, for $k\geq 1$, we have
$\mathfrak{c}_{nk}$ $=\langle\!\langle X_{n}-\mu,\phi_{k}\rangle\!\rangle=\frac{1}{\sqrt{l_{k}}}\sum_{j=1}^{N}\sqrt{\pi_{j}}[v_{k}]_{j}\langle\!\langle X_{n}-\mu,X_{j}-\mu\rangle\!\rangle$ (74)
$=\frac{1}{\sqrt{l_{k}\pi_{n}}}\sum_{j=1}^{N}[v_{k}]_{j}\mathbf{M}_{nj}=\sqrt{\frac{l_{k}}{\pi_{n}}}[v_{k}]_{n}.$ (75)
Concerning the expansion of the data into the basis of function $\Psi$, we write
$\mathbf{M}=\left(\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}\left(\mathrm{I}_{\!N}-\mathbf{1}_{\!N}\Pi^{\top}\right)\mathbf{C}W^{1/2}\right)\left(\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}\left(\mathrm{I}_{\!N}-\mathbf{1}_{\!N}\Pi^{\top}\right)\mathbf{C}W^{1/2}\right)^{\top}.$ (77)
We note
$\mathbf{A}=\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}\left(\mathrm{I}_{\!N}-\mathbf{1}_{\!N}\Pi^{\top}\right)\mathbf{C}W^{1/2},$ (78)
such that $\mathbf{M}=\mathbf{A}\mathbf{A}^{\top}$. We also assume that $\phi_{1},\phi_{2},\dots$ the eigenfunctions of the covariance operator $\Gamma$ have a decomposition into the basis $\Psi$
$\phi_{k}(\cdot)=\begin{pmatrix}\phi_{k}^{(1)}(\cdot)\\ \vdots\\ \phi_{k}^{(P)}(\cdot)\end{pmatrix}=\begin{pmatrix}\psi^{(1)\top}(\cdot)b_{1k}\\ \vdots\\ \psi^{(P)\top}(\cdot)b_{Pk}\end{pmatrix},\quad\text{where}\quad b_{pk}=\left(b_{pk1},\dots,b_{pkK_{p}}\right)^{\top}.$ (79)
We have, for $p=1,\dots,P$,
$\left(\Gamma\phi_{k}\right)^{(p)}(\cdot)$ $=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}C_{p,q}(\cdot,s_{q})\phi_{k}^{(q)}(s_{q}){\rm d}s_{q}$
$=\sum_{q=1}^{P}\int_{\mathcal{T}_{q}}\Psi(\cdot)^{(p)\top}\mathbf{C}^{(p)\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}^{(q)}\Psi^{(q)}(s_{q})\Psi^{(q)}(s_{q})^{\top}b_{qk}{\rm d}s_{q}$
$=\Psi(\cdot)^{(p)\top}\mathbf{C}^{(p)\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\sum_{q=1}^{P}\mathbf{C}^{(q)}\int_{\mathcal{T}_{q}}\Psi^{(q)}(s_{q})\Psi(s_{q})^{(q)\top}{\rm d}s_{q}b_{qk}$
$=\Psi(\cdot)^{(p)\top}\mathbf{C}^{(p)\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\sum_{q=1}^{P}\mathbf{C}^{(q)}\mathbf{W}^{(q)}b_{qk}.$
This equation is true for all $p=1,\cdots,P$, this can be rewritten with matrices as
$\Gamma\phi_{k}(\cdot)=\Psi(\cdot)^{\top}\mathbf{C}^{\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}\mathbf{W}b_{k}.$ (80)
From the eigenequation, we have that
$\Gamma\phi_{k}(\cdot)=\lambda_{k}\phi_{k}(\cdot)\Longleftrightarrow\Psi(\cdot)^{\top}\mathbf{C}^{\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}\mathbf{W}b_{k}=\lambda_{k}\Psi(\cdot)^{\top}b_{k}.$ (81)
Since this equation must be true for all $t_{p}\in\mathcal{T}_{p}$, this imply the equation
$\mathbf{C}^{\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}\mathbf{W}b_{k}=\lambda_{k}b_{k}.$ (82)
As the eigenfunctions are assumed to be normalized, $\left|\!\left|\!\left|\phi_{k}\right|\!\right|\!\right|^{2}=1$. And so, $b_{k}^{\top}Wb_{k}=1$. Let $u_{k}=W^{1/2}b_{k}$. Then, from Equation (82), we obtain
$\mathbf{W}^{1/2}\mathbf{C}^{\top}\left(\text{diag}\{\pi_{1},\dots,\pi_{N}\}-\Pi\Pi^{\top}\right)\mathbf{C}\mathbf{W}^{1/2}u_{k}=\lambda_{k}u_{k}\Longleftrightarrow\mathbf{A}^{\top}\mathbf{A}u_{k}=N\lambda_{k}u_{k}.$ (83)
From the eigendecomposition of the matrix $M$, we get
$\mathbf{M}v_{k}=l_{k}v_{k}\Longleftrightarrow\mathbf{A}\mathbf{A}^{\top}v_{k}=l_{k}v_{k}.$ (84)
The equations (83) and (84) are eigenequations in the classical PCA case, with the duality $X^{\top}X$ and $XX^{\top}$. Following Pagès, (2014); Härdle and Simar, (2019), we find that, for $1\leq k\leq K$,
$\lambda_{k}=l_{k},\quad v_{k}=\frac{1}{\sqrt{l_{k}}}\mathbf{A}u_{k}\quad\text{and}\quad u_{k}=\frac{1}{\sqrt{l_{k}}}\mathbf{A}^{\top}v_{k}.$ (85)
And finally, to get the coefficient of the eigenfunctions, for $1\leq k\leq K$,
$b_{k}=\mathbf{W}^{-1/2}u_{k}=\frac{1}{\sqrt{l_{k}}}\mathbf{C}^{\top}\left(\mathrm{I}_{\!N}-\Pi\mathbf{1}_{\!N}^{\top}\right)\text{diag}\{\sqrt{\pi_{1}},\dots,\sqrt{\pi_{N}}\}v_{k}.$ (86)

### v2

@@ B.1 General framework
Therefore, in order to have an orthonormal basis of eigenfunctions, we normalise the eigenfunctions $\phi_{k}$ from Equation (A.28) by $1/\sqrt{l_{k}}$. Concerning the estimation of the scores, for $n=1,\dots,N$, for $k\geq 1$, we have
$\mathfrak{c}_{nk}$ $=\langle\!\langle X_{n}-\mu,\phi_{k}\rangle\!\rangle=\frac{1}{\sqrt{l_{k}}}\sum_{j=1}^{N}\sqrt{\pi_{j}}[\boldsymbol{u}_{k}]_{j}\langle\!\langle X_{n}-\mu,X_{j}-\mu\rangle\!\rangle$ (A.29)
$=\frac{1}{\sqrt{l_{k}\pi_{n}}}\sum_{j=1}^{N}[\boldsymbol{u}_{k}]_{j}\mathbf{M}_{nj}=\sqrt{\frac{l_{k}}{\pi_{n}}}[\boldsymbol{u}_{k}]_{n}.$ (A.30)
If we assume that the observations are equally weighted, i.e., $\pi_{n}=1/N,n=1,\dots,N$, we get the equalities (36), (37) and (38).
Appendix C More results
C.1 Simulation
We present the results of the simulations in the sparse and noisy cases. Figures 11, 13 and 15 present the boxplots of the RSE, ISE and MRSE, respectively, for the noisy case. To generate noisy data, we consider the model (11) where $\sigma^{2}=0.25$. For the (Tensor) PCA method, we first smooth the 2-dimensional data using P-Splines smoothing and estimate a smooth version of the mean and covariance functions for the one-dimensional data using P-Splines smoothing. For the 2D/1D B-Splines and Gram methods, all the observations have been smoothed using P-Splines smoothing beforehand. In every case, the penalty involved in P-Splines smoothing has been estimated using cross-validation.
Figure 11: RSE for the estimated eigenvalues for each method in the noisy case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
Figure 12: RSE for the estimated eigenvalues for each method in the sparse case. We set $N=250$, $M^{(1)}=101\times 51$ and $M^{(2)}=201$. For the case of medium sparsity, we remove $50\%-70\%$ of the sampling points and for the case of high sparsity, we remove $90\%-95\%$ of the sampling points.
Figures 12, 14 and 16 present the boxplots of the RSE, ISE and MRSE, respectively, for the sparse case. In this case, we only consider $N=250$, $M^{(1)}=101\times 51$, $M^{(2)}=201$ and no noise. We consider medium sparsity, where $50\%-70\%$ of the sampling points have been removed and high sparsity, where $90\%-95\%$ of the sampling points have been removed. For the (Tensor) PCA method, we first interpolate the 2-dimensional data to have the observations reguarlarly sampled on a grid and estimate a smooth version of the mean and covariance functions for the one-dimensional data using P-Splines smoothing. For the 2D/1D B-Splines method, all the observations have been smoothed using P-splines smoothing with a penalty estimated by cross-validation. For the Gram method, the observations have been linearly interpolated to estimate the inner-product matrix.
Figure 13: ISE for the estimated eigenfunctions for each method in the noisy case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
Figure 14: ISE for the estimated eigenfunctions for each method in the sparse case. We set $N=250$, $M^{(1)}=101\times 51$ and $M^{(2)}=201$. For the case of medium sparsity, we remove $50\%-70\%$ of the sampling points and for the case of high sparsity, we remove $90\%-95\%$ of the sampling points.
Figure 15: MRSE for the reconstructed curves for each method in the noisy case. $N$ is the number of observations, $M$ is the number of sampling points per curve (the first two numbers are for the images and the last one is for the curves).
Figure 16: MRSE for the reconstructed curves for each method in the sparse case. We set $N=250$, $M^{(1)}=101\times 51$ and $M^{(2)}=201$. For the case of medium sparsity, we remove $50\%-70\%$ of the sampling points and for the case of high sparsity, we remove $90\%-95\%$ of the sampling points.
C.2 Application
Here, we present the estimation of the eigencomponents of the shooting data using the (Tensor) PCA (see Figure 17) and 2D/1D B-Splines (see Figure 18) methods. The results are similar to those found using the Gram method for the eigenfunctions and the percentage of variance explained for each eigenfunctions.
Figure 17: The estimated mean surfaces (first column) and the estimated eigenfunctions (second to fifth columns) for the shots dataset using the (Tensor) PCA method.
Figure 18: The estimated mean surfaces (first column) and the estimated eigenfunctions (second to fifth columns) for the shots dataset using the 2D/1D B-splines method.